Grade 12th passAlgebra

10 digits numbers are formed such that all the digits 0 to 9 must be used (e.g.1234567890) and they are also divisible by 11111. The digit in 10th place in smallest of such number and the digit at unit place in greatest of such number.........

Ritik
7 Years agoGrade 12th pass

1 Answer

Saurabh Koranglekar
6 Years ago
We know that
11111 × 90000 = 999990000
By intuition, we can conclude that 10000 is the smallest number which when added to
999990000 will give a 10 digited sum.
As 11111 > 10000,
11111*90001will be the smallest 10 digit number satisfying a given condition.
Needless to tell,
its 10's digit =1.

By using a bit common sense, we can predict that the largest 10 digit multiple of 11111 should be9999999999.
Unit digit =9.

Again let's use a bit common sense instead of using horrible A. P. formulae.
Before directly jumping to our problem, I'd love to give some introduction.
Suppose we've to find the multiples of 5 between 28 & 52.
What shall we do ?
The smallest multiple = 30.
The greatest multiple = 50.
30/5 = 6 & 50/5 = 10.
Total number of multiples =
(10-6) + 1 =5.
Now, let's turn towards our problem.
Now, I think the next steps are very clear.
We have :
11111*90001/11111 =90,001.
9999999999/11111 =9,00,009.
Number of 10 digit multiples of 11111 =9,00,009 - 90,001 + 1
= 9,00,009 - 90,000
=8,10,009.